3xy + y = 15
Solve for Y

Answer by mananth(16946)

3xy + y = 15
Solve for Y
..
y(3x+1)=15
divide by (3x+1)
y(3x+1)/(3x+1) = 15/(3x+1)
y= 15/(3x+1)